# Variation in the size of the Photon Sphere and Black Hole Shadow in the Modified Gravity

### v2

@@ III.2 Shadow in the Kottler space-time for a static observer
Here we discussed the effect of the cosmological constant on BHS and photon sphere. We considered the Kottler metric Kottler:1918A ; Perlick2021CalculatingBH
@@
Here we find two horizon Wang:2022kvg ; Perlick2021CalculatingBH for $0<\Lambda<(3m)^{-2}$ and also find out that the ratio of impact parameter and observer distance does not equal to angular radius of the black hole shadow for static observer just like in Schwarzschild space-time. Therefore we considered commoving observer with cosmological accelerated expansion.
@@ III.3 Shadow of Kerr black hole
Kerr black hole is an uncharged black hole with rotational motion. We have discussed the space-time around the black hole in vacuum with the help of Kerr metric Kerr:1963ud ; Boyer:1966qh ; Perlick2021CalculatingBH is
@@
Using above parameters, we have calculated the coordinates$(x,y)$ of a point in the observer plane and static observer is locally at $(r_{o},\theta_{o})$ in zero-angular momentum frame of reference 1973Bardeen ; Wang:2022kvg ; Johannsen:2013vgc . If the observer is at infinity $(r_{o}\rightarrow\infty)$ means far away from the black hole. We have Wang:2022kvg
@@
where -- and $+$ denote prograde and retrograde respectively. The deviation parameter is
$\delta=\frac{D}{R_{sh}}$ (59)
@@
$D=\frac{x_{min}+x_{max}}{2}$ (60)
where $x_{max}$ and $x_{min}$ are the maximum and minimum abscissae, respectively of the ring, respectively This is the deviation of black hole shadow from circle. The average radius of the photon ring is Johannsen:2013vgc
[…]
@@ IV Black hole shadow and Photon sphere in $f(R)$ gravity
The modified theory of gravity explain unsolved problems elegantly Sharma:2022fiw ; Sharma:2022tce ; Sharma:2019yix ; KumarSharma:2022qdf . In this article, we have used $f(R)$ background in place of GR background for the black hole shadow and photon sphere. In $f(R)$ gravity, modifications to the spacetime curvature can alter trajectories of photon, changing the size and shape of the shadow nojiri2024black ; nojiri2025black . These deviations appear because of the specific form of $f(R)$, the black hole’s spin, and surrounding matter fields.
@@ IV.1 Field equations in $f(R)$ gravity
We considered static and spherical symmetric spacetime metric capozziello2019maximum given by equation (18), where
@@
$G_{eff}=G_{N}/F$, and $F=f^{\prime}(R)=df(R)/dR$. Einstein-Hilbert action in $f(R)$ gravity is de:2010f ; nojiri:2017modified
@@
For spherically symmetric case, we used constant curvature solution of $f(R)$ model i.e. $R=R_{dS}=constant$. The equation (62) becomes capozziello2019maximum
@@ V Conclusion
In this paper, we have found that the size of the black hole photon sphere and shadow changed with metric tensor and new scalar degree fo freedom $F$ appears in the modified theory of gravity gravity. Indeed, There is no need to introduced two components separately in the modified gravity. Therefore, we preferred $f(R)$ theory of gravity. We obtained generalized formulae of photon sphere 78 and black hole shadow 85 for $f(R)$ theory.
In Kottler spacetime, there is two horizon and we can also find two horizon in $f(R)$ gravity from the equation 85. If we considered the value $F=1$ and $R_{dS}=4\Lambda$, we got similar expression of black hole shadow in Kottler spacetime. That means, we recovered the expression in GR background.

### v3

@@ III.2 Shadow in the Kottler space-time for a static observer
Here we discussed the effect of the cosmological constant on BHS and photon sphere. We considered the Kottler metric Kottler (1918); Perlick and Tsupko (2021)
@@
Here we find two horizon Wang et al. (2022b); Perlick and Tsupko (2021) for $0<\Lambda<(3m)^{-2}$ and also find out that the ratio of impact parameter and observer distance does not equal to angular radius of the black hole shadow for static observer just like in Schwarzschild space-time. Therefore we considered commoving observer with cosmological accelerated expansion.
@@ III.3 Shadow of Kerr black hole
Kerr black hole is an uncharged black hole with rotational motion. We have discussed the space-time around the black hole in vacuum with the help of Kerr metric Kerr (1963); Boyer and Lindquist (1967); Perlick and Tsupko (2021) is
@@
Using above parameters, we have calculated the coordinates(x,y)(x,y) of a point in the observer plane and static observer is locally at $(r_{o},\theta_{o})$ in zero-angular momentum frame of reference Bardeen (1973); Wang et al. (2022b); Johannsen (2013). If the observer is at infinity $(r_{o}\rightarrow\infty)$ means far away from the black hole. We have Wang et al. (2022b)
@@
where - and ++ denote prograde and retrograde respectively. The deviation parameter is
$\delta=\frac{D}{R_{sh}},$ (59)
@@
$D=\frac{x_{min}+x_{max}}{2},$ (60)
where $x_{max}$ and $x_{min}$ are the maximum and minimum abscissae, respectively of the ring, respectively This is the deviation of black hole shadow from circle. The average radius of the photon ring is Johannsen (2013)
$\langle\bar{R}\rangle=\frac{1}{2\pi}\int_{0}^{2\pi}\bar{R}d\alpha,$ (61)
@@ IV Black hole shadow and Photon sphere in $f(R)$ gravity
The modified theory of gravity explain unsolved problems elegantly Sharma and Verma (2022a, c); Sharma et al. (2020b); Kumar Sharma and Verma (2022). In this article, we have used $f(R)$ background in place of GR background for the black hole shadow and photon sphere. In $f(R)$ gravity, modifications to the spacetime curvature can alter trajectories of photon, changing the size and shape of the shadow Nojiri and Odintsov (2024, 2025). These deviations appear because of the specific form of $f(R)$, the black hole’s spin, and surrounding matter fields.
@@ IV.1 Field equations in $f(R)$ gravity
We considered static and spherical symmetric spacetime metric Capozziello et al. (2019) given by equation (18),
@@
where $G_{eff}=G_{N}/F$, and $F=f^{\prime}(R)=df(R)/dR$. Einstein-Hilbert action in $f(R)$ gravity is De Felice and Tsujikawa (2010); Nojiri et al. (2017)
@@
For spherically symmetric case, we used constant curvature solution of $f(R)$ model i.e. $R=R_{dS}=constant$. The equation (62) becomes Capozziello et al. (2019)
@@ V Conclusion
In this paper, we have found that the size of the black hole photon sphere and shadow changed with metric tensor and new scalar degree fo freedom $F$ appears in the modified theory of gravity gravity. Indeed, There is no need to introduced two components separately in the modified gravity. Therefore, we preferred $f(R)$ theory of gravity. We obtained generalized formulae of photon sphere (78) and black hole shadow (85) for $f(R)$ theory.
In Kottler spacetime, there is two horizon and we can also find two horizon in $f(R)$ gravity from the equation (85). If we considered the value $F=1$ and $R_{dS}=4\Lambda$, we got similar expression of black hole shadow in Kottler spacetime. That means, we recovered the expression in GR background.